Abdullah has negotiated a price of $25,800 for a new SUV. Now he must choose between 0% financing for 48 months or a $3,000 rebate. If he chooses the rebate he can obtain a loan for the balance at 4.5% compounded monthly for 48 months at a local bank. Which option should he choose?
Answer by rfer(16322)

He will save $1487.37 by taking the no interest loan.
25800-3000=22800(1+.045/12)^48
A=22800(1.19681438)
A=$27287.37
27287.37-25800.00=$1487.37